Current and Quick Ratios The Nelson Company has $800,000 in current assets and $400,000 in current liabilities. Its initial inventory level is $240,000, and it will raise funds as additional notes payable and use them to increase inventory.

1- How much can Nelson's short-term debt (notes payable) increase without pushing its current ratio below 1.5? Round your answer to the nearest cent.

2- What will be the firm's quick ratio after Nelson has raised the maximum amount of short-term funds? Round your answer to two decimal places.
Answer #1

1)current ratio = (current assets+ increase in inventory)/(currentliabilities+ increase in short term debt)

note :increase in inventory= increase in short term debt

1.5 = (800000+x)/(400000+x)

x = 400000 = increase in notes payable

2)

Quick or acid test ratio = (current assets-inventory)/current liablities

=(800000-240000)/(400000+400000) = 0.7